# Constraining Participation: Affordances of Feedback Features in Interfaces to Large Language Models

### v1

@@ 3. Method
We conducted an online survey to gather insights from ChatGPT users who had submitted feedback using the features in the browser-based interface. Through the survey, we aimed to explore feedback practices and perspectives of ChatGPT users to understand which users submit feedback, how they submit feedback, and why they submit feedback. We then drew on the insights from the survey to analyse the feedback features through the mechanisms and conditions framework of affordances. Below, we detail the process we used to conduct the survey and the analysis, and we then describe the limitations of our approach. In the following section, we present the insights from our analysis.
@@ 3.1.1. Pilot Survey
We conducted two pilot surveys in March 2023 with 32 respondents from our university to refine our survey instrument’s clarity, flow, and length. In late March 2023, the Human Ethics Review Committee at our university approved our recruitment plan and final survey instrument.
@@ 3.1.2. Survey Instrument
The survey instrument consisted of 45 questions—approximately 21 were consistent across respondents, and the survey logic determined the remainder. The survey asked respondents about their frequency and types of ChatGPT usage, experiences with submitting feedback through the ChatGPT interface and external channels, reasons for submitting or not submitting feedback, perceptions of the feedback process, the likelihood of providing future feedback through the interface and external channels, and demographic characteristics (including the AI, ML, and data science background of respondents). Most questions were multiple-choice, with a few open-ended questions to gather more detailed insights and reflections from respondents.
@@ 3.1.3. Recruitment and Sampling
We administered the public survey on Qualtrics in May 2023 and used a multi-pronged approach to recruit participants. Initially, the authors posted the survey on their social media accounts (Twitter, Mastodon, and LinkedIn), forums frequented by active ChatGPT users (GitHub forums, Hugging Face forums, the OpenAI community forum, and Kaggle forums), and on the r/ChatGPT Discord server.
We also recruited participants through Prolific to address non-response bias and expand the respondent pool beyond those engaged with AI, ML, or data science in research or development capacities. We applied screening strategies on Prolific to invite relevant respondents to the final survey. Firstly, we used Prolific filters to identify respondents from any country, reporting fluency in English, an approval rate greater than or equal to 95%, and 10 or more previous submissions. Secondly, we invited filtered respondents to a custom screening survey with two questions asking participants about chatbot usage and feedback submission. [^1] ChatGPT usage was not an option for filters on Prolific in May 2023. Finally, we invited respondents approved after the screening survey to the final survey at three separate intervals to cover working hours in regions with different time zones. Respondents to the final survey on Prolific received the equivalent of £9/hr for their time.
@@ 3.1.4. Respondents
The survey received responses from 526 individuals—after removing incomplete surveys, failed attention checks [^2] Only Prolific surveys., and screened surveys [^3] Valid IP Address, “Yes” to “Read and agree to Participant Information Sheet”, “Yes” to “Has used ChatGPT”.—with a median completion time of five minutes and 11 seconds. The respondents were predominantly young, with over three quarters aged 18–34 (42.4% aged 18–24, 35.1% aged 25–34, n=524). The gender distribution was skewed towards men (63.1%), with 34.9% identifying as women, 1.2% as non-binary or third gender, and 0.8% preferring not to say or preferring to self-describe (n=521). Geographically, the respondents were primarily from Europe (55.4%), followed by Africa (15.6%), the United Kingdom (10.3%), and Oceania (6.5%) (n=507). The most common languages spoken by respondents were English (UK, US, or AU) (61.1%), Portuguese (13.4%), and Polish (12.2%) (n=509). [^4] Respondents could select up to three languages they speak daily.
Respondents reported varying levels of engagement with AI-related topics (e.g., through news articles, research papers, podcasts, etc.), with 63.1% at least weekly, 18.9% monthly, and only 18.0% rarely or never (n=523). A substantial proportion (41.0%) had used AI, ML, or data science tools beyond basic interaction or exploration before using ChatGPT (n=522). Furthermore, 26.5% had completed formal education or training in these areas (n=520), and 21.3% were working or had previously worked in a related field (n=521).
@@ 3.2. Data Analysis
Initially, we performed frequency analysis for most questions (see Appendix A). For certain variables measured on a 5-point Likert scale, we transformed the responses into three categories (e.g., Agree, Neutral, Disagree; or Likely, Neutral, Unlikely) to simplify the analysis and interpretation. We then performed chi-square tests of independence ($\chi^{2}$) to analyse relationships between responses, using Cramér’s V to determine the strength of the associations (Cohen, 2013). Before these analyses, we removed rows with missing values for demographics and experience factors and grouped categories with low frequencies for relevant questions (e.g., consolidating age groups and geographic regions). To address the multiple comparisons problem, we applied the Benjamini-Hochberg procedure to control the false discovery rate across all tests simultaneously (Benjamini and Hochberg, 1995). We report statistically significant results based on the adjusted $p$-values (adj. $p$-value). All statistical tests were conducted at a 95% confidence level (df=degrees of freedom; see Appendix B for full results).
@@ 3.3. Limitations
Our methodology has several important limitations. Firstly, we recognise that our survey dataset does not statistically represent any group—neither ChatGPT users in general nor the subset of ChatGPT users who have submitted feedback through the interface (Tufekci, 2014). We sampled a subset of ChatGPT users in May 2023, though the networks of the authors and the population of users on Prolific influence our sample. The number and characteristics of ChatGPT users in general are not freely available, nor are the number or characteristics of ChatGPT users who have submitted feedback. This survey instrument was conducted to sample that population within a constrained budget. As such, our survey insights are exploratory and not generalisable to the population of ChatGPT users. Nevertheless, these insights are a useful starting point for understanding user practices and perspectives related to feedback features in interfaces to LLMs.
Secondly, our survey relies on self-reported data from respondents. We did not create a primary dataset of feedback instances, as we could not access a dataset of feedback instances submitted through the ChatGPT interface. Instead, we collected a secondary dataset of the type of feedback (e.g., thumbs-up/down, open-text feedback, etc.) and the broad categories of feedback (e.g., correcting technical errors/bugs, social and ethical information, etc.) provided by respondents. There may be discrepancies between what users reported and their actual feedback behaviour. We note, however, that the survey was conducted primarily to understand the practices and perspectives of ChatGPT users who have submitted feedback rather than to analyse the content of the feedback itself.
@@
Finally, while survey findings provide useful information about user practices and perspectives concerning feedback, there are limitations to using this data for our affordances analysis. The survey provides a snapshot of user perspectives and experiences at a point in time. The ChatGPT interface and feedback features have and will continue to change, so the affordances of those feedback features and user behaviours may change accordingly. However, as of May 2024, the basic functionality has not changed significantly, though users are prompted with feedback categories where they were previously provided with an open-text box (see Fig. 5), and a user may be requested to submit pairwise feedback (see, e.g., Fig. 6). We also note that user interviews or observations could provide more in-depth information about how users perceive and interact with the feedback features, though we argue that survey data is sufficient for our high-level analysis through the mechanisms and conditions framework of affordances.
Figure 5. Additional feedback categories after selecting thumbs-down - ChatGPT (May 2024)
[…]
An image of the ChatGPT interface showing a feedback window that appears after a user selects the thumbs-down option. The window is titled “Provide additional feedback” and presents multiple categorised feedback options for users to specify why they disliked the response. Categories include “Don’t like the style”, “Not factually correct”, “Didn’t follow instructions”, “Refused when it shouldn’t have”, “Being lazy”, “Unsafe or problematic”, and “Other”. At the bottom of the window is a single-line text box labelled “(Optional) Feel free to add specific details” for users to provide more specific feedback.
Figure 6. Pairwise feedback options - ChatGPT (May 2024, content of responses hidden)
[…]
An image of the ChatGPT interface displaying a pairwise comparison feature. At the top of the window, the text reads “Which response do you prefer? Your answer will help make ChatGPT better”. Below, the image shows two text areas side by side containing different generated responses, though the content of the responses is hidden for privacy. This interface design facilitates direct comparison between two alternative responses.
@@ 4. Insights and Analysis
This section reports insights from the survey regarding ChatGPT usage, feedback practices, and respondents’ motivations for submitting feedback. Throughout the section, we draw on the insights to inform our analysis of the affordances of feedback features in the ChatGPT interface.
Our analysis suggests how the design of feedback features imposes constraints on end-user participation. In summary, the interface requests feedback formatted in the ‘helpful, honest, and harmless’ framework, demands interaction before allowing feedback submission, and encourages simple, frequent, individualised, and unidirectional feedback, particularly thumbs-up/down ratings and feedback after regenerating a response. However, the interface discourages collective feedback and refuses feedback modification over time. Additionally, our survey indicates that the feedback features may appeal more to users with frequent engagement, prior experience, and knowledge of AI, ML or data science.
@@ 4.1. Mechanisms of Affordance
Our survey reveals how users interact with ChatGPT and its feedback features, which influence the type and quality of data fed back into RLHF workflows. Respondents used ChatGPT frequently and in diverse ways. Over 71.6% (n=525) used ChatGPT at least weekly, and the most common use cases were answering specific questions (e.g., factual, technical, or theoretical questions), generating text (e.g., drafting emails, writing essays or reports), and academic research assistance (e.g., generating citations, summarising articles).
@@
This usage pattern is significant because a person cannot directly influence ML workflows outside that system. A user must prompt the dialogue agent before they can submit feedback. This design feature effectively demands interaction with an LLM before allowing participation through feedback submission, establishing the first key mechanism of affordance we observed.
4.1.1. Submitting Feedback
The feedback features in the ChatGPT interface are designed to elicit individual responses, which effectively discourages collective feedback or discussions among users. This individualisation extends to the permanence of feedback: users cannot modify their input once submitted, even if they gain new information or their perspective changes. While users could theoretically input the same prompt a second time and submit feedback, the stochastic nature of LLMs means each response is likely to be unique, making direct comparisons or revisions challenging. This design choice reinforces the individual nature of the feedback process and limits opportunities for evolving, collective input.
@@ 4.1.1. Submitting Feedback
Our survey findings indicate that the ChatGPT interface encourages thumbs-up/down feedback and feedback after regenerating a response. Most respondents (84.2%, n=526) indicated they had submitted feedback through one of the three methods in the ChatGPT interface. Thumbs-up/down feedback was the most common (76.0%, n=526), followed by feedback after regenerating responses (62.9%, n=526), while open-text feedback was less common (29.0%, n=526).
Respondents submitting thumbs-up/down feedback also reported doing so most frequently (53.3% every, most, or about half the times they had used ChatGPT, n=400), followed by feedback after regenerating a response (47.4% every, most, or about half the times they had used ChatGPT, n=331). Respondents submitting open-text feedback reported doing so least frequently (26.4% every, most, or about half the times they had used ChatGPT, n=152).
@@
Respondents who submitted open-text feedback (n=152) were primarily motivated by improving the performance and capabilities of the system. We asked this group of respondents to reflect on the most recent time they submitted open-text feedback and report their motivations for doing so. Respondents could select multiple motivations, but those who selected multiple motivations were forced to choose their primary motivation in the following question. Taking the weighted frequencies, the primary motivations were (in order) to: “contribute to AI development”, “comment on use-case specific feedback”, “report a technical issue or bug”, and “suggest new features or improvements” (see Table 12 in Appendix A). Expressing social or ethical concerns about system outputs was the least common motivation (excluding the range of individual responses in the “other” option).
4.1.2. Reflecting on Feedback
@@
We also asked respondents who submitted open-text feedback what feedback they thought was most useful for improving ChatGPT’s performance or impact. The responses revealed a hierarchy of feedback themes. The most prominent themes were technical issues (bugs, glitches, incorrect information) and system performance (efficiency, speed, and quality of responses). Less common but still significant were concerns about accuracy (‘hallucinations’ or false information) and the reliability and truthfulness of responses. Though less frequent again, some users highlighted social and ethical concerns (including harmful, biased, or misleading responses). Finally, some feedback focused on suggesting new features or improvements.
Most respondents were ambivalent about whether their reasons for submitting feedback had been addressed—47.4% neither agreed nor disagreed, 23.7% somewhat agreed, and 16.4% somewhat disagreed (n=152). Few of those who submitted open-text feedback indicated they had received follow-up communication or updates from OpenAI (11.2%, n=152), though the vast majority of those who had indicated they were satisfied with the response or resolution.
@@
The feedback features in the ChatGPT interface allow for open-text feedback, including commentary on social and ethical issues. However, our survey results reveal that users predominantly focus on performance-related concerns when submitting feedback. This focus appears to be shaped by the affordances of the feedback system itself rather than by explicitly discouraging social or ethical feedback. Despite OpenAI’s initial blog post requesting feedback on risks and harms, users who responded to our survey seem to have interpreted the purpose of their participation primarily through the lens of improving ChatGPT’s performance. This pattern highlights how the design of feedback features can effectively constrain user perspectives and shape the nature of user input.

### v2

@@ 3. Method
We conducted an online survey to gather insights from ChatGPT users who had submitted feedback using the features in the browser-based interface. Through the survey, we aimed at exploring feedback practices and perspectives of ChatGPT users to understand which users submit feedback, how they submit feedback, and why they submit feedback. We then drew on the insights from the survey to analyse the feedback features through the mechanisms and conditions framework of affordances. Below, we detail the process we used to conduct the survey and the analysis, and we then describe the limitations of our approach. In the following section, we present the insights from our analysis.
@@ 3.1.1. Pilot Survey
We conducted two pilot surveys in March 2023 with 32 respondents from the Australian National University (ANU) to refine our survey instrument’s clarity, flow, and length. In late March 2023, the ANU approved our recruitment plan and final survey instrument.
@@ 3.1.2. Survey Instrument
The survey instrument consisted of 36 substantive questions exploring respondents’ frequency and types of ChatGPT usage, experiences submitting feedback through the ChatGPT interface and external channels, reasons for submitting or not submitting feedback, perceptions of the feedback process, the likelihood of providing future feedback, and demographic characteristics (see Appendix A for survey questions). Most questions were multiple-choice, with several open-ended questions included to gather more detailed insights and reflections.
@@ 3.1.3. Recruitment and Sampling
We administered the public survey on Qualtrics in May 2023 and used a multi-pronged approach to recruit a diverse range of participants. This approach was necessary as the ChatGPT user population was not well-defined at the time of our study. Initially, we posted the survey on our social media accounts (X (formerly known as Twitter), Mastodon, and LinkedIn), forums frequented by active ChatGPT users (GitHub forums, Hugging Face forums, the OpenAI community forum, and Kaggle forums), and on the r/ChatGPT Discord server.
We then expanded recruitment through Prolific to address potential non-response bias and broaden the respondent pool. We applied screening strategies on Prolific to invite suitable respondents to the final survey. Firstly, we used Prolific filters to identify respondents from any country, reporting fluency in English, an approval rate greater than or equal to 95%, and 10 or more previous submissions. Secondly, we invited filtered respondents to a custom screening survey with two questions asking participants about chatbot usage and feedback submission. [^1] ChatGPT usage was not an option for filters on Prolific in May 2023. Finally, we invited respondents approved after the screening survey to the final survey at three separate intervals to cover working hours in regions with different time zones. Respondents to the final survey on Prolific received the equivalent of £9/hr for their time.
@@ 3.1.4. Respondents
The survey received responses from 526 individuals—after removing incomplete surveys, failed attention checks, [^2] Only Prolific surveys. and screened surveys [^3] Valid IP Address, “Yes” to “Read and agree to Participant Information Sheet”, “Yes” to “Has used ChatGPT”.—with a median completion time of five minutes and 11 seconds. We recruited 82 respondents through the first recruitment step and 445 through the second. As shown in Table 1, our sample includes perspectives from users with varying levels of technical experience and engagement with AI-related topics (e.g., through news articles, research papers, podcasts). Most respondents did not have formal AI training (73.5%) or relevant work experience (78.7%), suggesting our recruitment approach reached beyond technically experienced users. However, respondents were predominantly young, with over three quarters aged 18–34, and the gender distribution was skewed toward men. Geographically, the majority of respondents were from Europe, and the most common languages spoken by respondents were English (UK, US, or AU), Portuguese, and Polish. [^4] Respondents could select up to three languages they speak daily.
Table 1. Respondent Demographics
Characteristic Distribution
Age ($n$ = 524) 18-34 (77.5%), 35-54 (19.6%), 55+ (2.3%)
Gender ($n$ = 521) Man (63.1%), Woman (34.9%), Non-binary/Third gender (1.2%), Prefer not to say/Prefer to self-describe (0.8%)
Region ($n$ = 507) Europe (55.4%), Africa (15.6%), UK (10.3%), Oceania (6.5%), Other (12.2%)
AI Experience Prior use of AI/ML tools (41.0%, $n$ = 522), Formal education in AI/ML (26.5%, $n$ = 520), Working in AI/ML field (21.3%, $n$ = 521)
AI Engagement ($n$ = 523) Daily/Weekly (63.1%), Monthly (18.9%), Rarely/Never (18.0%)
@@ 3.2. Data Analysis
Initially, we performed frequency analysis for most questions (see Tables 937). For certain variables measured on a five-point Likert scale, we transformed the responses into three categories (e.g., Agree, Neutral, Disagree; or Likely, Neutral, Unlikely) to enhance statistical validity for chi-squared tests (by increasing cell counts and reducing the risk of minimum frequency violations) and improve interpretability. To assess the potential impact of this categorisation on our ordinal Likert scale data we performed sensitivity analyses on key relationships using the original five-point scales. These analyses showed consistent significance patterns in most cases, with similar effect sizes across both approaches.
We then performed chi-squared tests of independence ($\chi^{2}$) to analyse relationships between responses, using Cramér’s V to determine the strength of the associations (Cohen, 2013). Before these analyses, we removed rows with missing values for demographics and experience factors and grouped categories with low frequencies for relevant questions (e.g., consolidating age groups and geographic regions). To address the multiple comparisons problem, we applied the Benjamini-Hochberg procedure to control the false discovery rate across all tests simultaneously (Benjamini and Hochberg, 1995). We report statistically significant results based on the adjusted $p$-values (adj. $p$-value). All statistical tests were conducted at a 95% confidence level (df=degrees of freedom; see Appendix C for full results).
@@ 3.3. Limitations
Our methodology has several important limitations. Firstly, the survey dataset does not statistically represent all ChatGPT users or the subset who submit feedback, limiting the generalisability of our findings (Tufekci, 2014). We sampled a subset of users in May 2023, influenced by the authors’ networks and the Prolific user population. Additionally, while necessary to engage with users during ChatGPT’s early release, sourcing some respondents from dedicated online forums might introduce a bias toward users inclined to discuss experiences publicly, potentially overlapping with those who share feedback through alternative channels. Our multi-pronged recruitment approach was pragmatic given that characteristics of the user base and feedback-submitting population were undefined at the time of study. Consequently, our survey insights are exploratory but useful as a starting point for understanding user practices and perspectives related to feedback features.
Additionally, our survey relies on self-reported data from respondents. We did not create a primary dataset of feedback instances, as we could not access a dataset of feedback instances submitted through the ChatGPT interface. Instead, we collected a secondary dataset of the type of feedback (e.g., thumbs-up/down, open-text feedback) and the broad categories of feedback (e.g., correcting technical errors/bugs, social and ethical information) provided by respondents. There may be discrepancies between what users reported and their actual feedback behaviour. We note, however, that the survey was conducted primarily to understand the practices and perspectives of ChatGPT users who have submitted feedback rather than to analyse the content of thefeedback itself.
@@
Finally, while survey findings provide useful information about user practices and perspectives concerning feedback, there are limitations to using this data for our affordances analysis. The survey provides a snapshot of user perspectives and experiences at a point in time. The ChatGPT interface and feedback features have and will continue to change, so the affordances of those feedback features and user behaviours may change accordingly. However, as of May 2024, the basic functionality has not changed significantly, though users are prompted with feedback categories where they were previously provided with an open-text box (see Figure 5), and a user may be requested to submit pairwise feedback (see, e.g., Figure 6). We also note that user interviews or observations could provide more in-depth information about how users perceive and interact with the feedback features, though we argue that survey data is sufficient for our high-level analysis through the mechanisms and conditions framework of affordances.
Figure 5. Additional feedback categories after selecting thumbs-down - ChatGPT (May 2024).:
Figure 6. Pairwise feedback options - ChatGPT (May 2024, content of responses hidden).:
@@ 4. Insights and Analysis
This section reports insights from the survey regarding ChatGPT usage, feedback practices, and respondents’ motivations for submitting feedback. Our analysis indicates how the design of feedback features imposes constraints on end-user participation, operating through various mechanisms and conditions that we explore below. First, we present findings about user interactions with ChatGPT, then examine each affordance mechanism, and finally discuss how these mechanisms operate through different conditions.
Respondents used ChatGPT frequently and in diverse ways. Over 71.6% (nn=525) used ChatGPT at least weekly, and the most common use cases were answering specific questions (e.g., factual, technical, or theoretical questions), generating text (e.g., drafting e-mails, writing essays or reports), and academic research assistance (e.g., generating citations, summarising articles).
@@ 4.1. Mechanisms of Affordance
4.1.1. Request and Demand
The ChatGPT interface requests feedback formatted in the “helpful, honest, and harmless” framework, as evident in the feedback submission options presented to users. This framing guides users toward specific types of evaluative feedback aligned with OpenAI’s stated development priorities.
@@ 4.1.1. Request and Demand
More fundamentally, the interface demands interaction with the system before allowing participation through one of the three modes for feedback submission. A user must prompt the dialogue agent before they can submit feedback. This design feature effectively makes system usage a prerequisite for participation in the feedback process, establishing a key mechanism of affordance we observed.
Table 2. Types of Feedback Submitted (nn=526)
Feedback Type Submitted Not Submitted
Thumbs-up/down feedback 76.0% 24.0%
Open-text feedback 29.0% 71.0%
Feedback after regenerating response 62.9% 37.1%
4.1.2. Encourage
@@
Our survey reveals several patterns in what forms of feedback the interface encourages. As shown in Table 2, a large majority of respondents (84.2%) submitted at least one form of feedback through the ChatGPT interface. Thumbs-up/down feedback was the most common, followed by feedback after regenerating responses, while open-text feedback was less common.
Furthermore, among those who submitted feedback, respondents submitting thumbs-up/down feedback reported doing so most frequently (28.8% every/most times), followed by feedback after regenerating a response (27.8% every/most times), while respondents submitting open-text feedback reported doing so least frequently (15.2% every/most times) (see Table 3). The observed preference for simpler feedback features may be reinforced by the interface design itself, where the thumbs-up/down feedback icons are visually prominent after each response, while accessing the open-text feedback requires an additional click.
Table 3. Frequency of Feedback Submission by Type
Frequency Thumbs-up/down Open-text After Regenerating
(nn=400) (nn=152) (nn=331)
Every/Most times 28.8% 15.2% 27.8%
About half the time 24.5% 11.2% 19.6%
Sometimes 34.8% 50.7% 36.3%
Rarely 12.0% 23.0% 16.3%
Total 100.0% 100.0% 100.0%
@@
The interface appears to particularly encourage performance-related feedback. Respondents who submitted open-text feedback were primarily motivated by improving the performance and capabilities of the system. We asked this group of respondents to reflect on the most recent time they submitted open-text feedback and report their motivations for doing so. Respondents could select multiple motivations, but those who selected multiple motivations were forced to choose their primary motivation in the following question. Table 4 shows the primary motivations in order, taking the weighted frequencies.
Table 4. Primary Motivations for Submitting Open-text Feedback (nn=152)
Motivation Weighted Frequency
To contribute to AI development 159
To comment on use-case specific feedback 94
To report a technical issue or bug 74
To suggest new features or improvements 73
To express social or ethical concerns about ChatGPT 32
Other 15
@@
Respondents were also asked what feedback they thought most useful for improving ChatGPT. Based on our thematic analysis, respondents emphasised technical issues (bugs, glitches) and system performance (efficiency, speed). Concerns about accuracy (“hallucinations” or false information) and the reliability or truthfulness of responses were less common, while social and ethical concerns (harmful, biased, or misleading responses) were mentioned even less frequently.
@@
In summary, while the feedback features in the ChatGPT interface allow for open-text feedback, including commentary on social and ethical issues, our respondents predominantly focused on performance-related concerns when submitting feedback. This focus appears to be shaped by the affordances of the feedback system itself rather than by explicitly discouraging social or ethical feedback. Despite OpenAI’s initial blog post requesting feedback on “harmful outputs” and “novel risks”, our survey respondents appear to have interpreted the purpose of their participation primarily through the lens of improving ChatGPT’s performance.
4.1.3. Discourage and Refuse
The feedback features in the ChatGPT interface are designed to elicit individual responses, which effectively discourages collective feedback or discussions among users. This individualisation extends to the permanence of feedback: the interface refuses modification of input once submitted, even if users gain new information or their perspective changes. While users could theoretically input the same prompt a second time and submit feedback, the stochastic nature of LLMs means each response is likely to be unique, making direct comparisons or revisions challenging. This design choice reinforces the individual nature of the feedback process and limits opportunities for evolving, collective input.
Most respondents were ambivalent about whether their reasons for submitting feedback had been addressed—47.4% neither agreed nor disagreed, 23.7% somewhat agreed, and 16.4% somewhat disagreed (nn=152). Few of those who submitted open-text feedback indicated they had received follow-up communication or updates from OpenAI (11.2%, nn=152), though the vast majority of those who had indicated they were satisfied with the response or resolution.
4.1.4. Allow
While the ChatGPT interface may constrain direct feedback, it allows for alternative channels of expression. Most respondents (80.4%, nn=526) did not indicate they shared feedback about ChatGPT through other channels or platforms. However, those who did primarily used X/Twitter (42%), followed by Facebook (18%) and Reddit (17%). Several respondents engaged with dedicated communities, including subreddits like r/ChatGPT and online forums like the OpenAI Developer Forum. Most respondents who indicated they had shared feedback through other channels or platforms (71.9%, nn=103) reported they were likely to continue to share feedback through those channels.